Which of the following functions are injective? Which are surjective?

a) f: Z → Z given by f(x) = x2 + 1.

b) g: N → N given by g(x) = 2x.

c) h: R → R given by h(x) = 5x - 1.


1
Expert's answer
2021-10-29T02:53:23-0400

a) "f:\ \mathbb{Z}\rightarrow \mathbb{Z} ,\quad f(x)=x^2+1"

1) "f(-1)=f(1)=2"

"f" is not injective

2) Since "f(x)=x^2+1\geq 1" , there doesn’t exists any "x\in\mathbb{Z}" such that "f(x)=0"

f is not surjective


b) "g:\ \mathbb{N}\rightarrow \mathbb{N}" , "g(x)=2^x"

1) If "g(x)=g(y)" , "2^x=2^y" , then "x=y"

f is injective

2) "g(x)=2^x=1" only if "x=0" , but "0\not \in \mathbb{N}"

f is not surjective


c) "h:\ \mathbb{R}\rightarrow \mathbb{R}" , "h(x)=5x-1"

1) If "h(x)=h(y)" , "5x-1=5y-1" , "5x=5y" , then "x=y"

f is injective

2) For all "y\in\mathbb{R}" there exists "x=\frac{y+1}{5}" such that "f(x)=5x-1=y+1-1=y"

f is surjective